# Complements of discriminants of real parabolic function singularities. II

### v4

@@ 2.5. The components for the $X_{9}^{1}$ singularities
Each polynomial (15) is invariant under the transformation $(x,y)\mapsto(y,x)$. The transformation $(x,y)\mapsto(-x,-y)$ maps each such polynomial to the analogous polynomial with the opposite value of parameter $t$. In particular, it permutes the connected components described by pictures $(d)$ and $(e)$.
@@ 2.6. Components for the $X_{9}^{2}$ singularities
The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the complement of the discriminant in some way. All these components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
@@
Each polynomial of degree four in ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by an affine transformation of ${\mathbb{R}}^{2}$ to the form
@@
Proof of Lemma 23. This lemma follows from Proposition 22 of [28] by the substitution $\tilde{x}=|a|^{\frac{-3}{4}}X$, $\tilde{y}=sign(a)|a|^{\frac{1}{4}}Y$, $t=|a|^{\frac{-5}{4}}$. $\Box$
@@ 3. Components of discriminant complements for $J_{10}$ singularities
Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (31) and (32), respectively. The formal graph and its virtual components for $J_{10}$ singularities are defined in the same way as for the case $X_{9}$, except that the frames of vanishing cycles consist of ten elements.
@@
In the following pictures, it is assumed that the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
@@ 3.2. Components for $J_{10}^{1}$ singularities
The following is an invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(e)$ from its mirror image f(x,y)f(x,-y). The four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement. Let us take an arbitrary point inside each of these four ovals. There is a single curve with the equation $x=ay^{3}+by^{2}+cy+d$ that passes through these four points. The coefficient $a$ of this equation cannot be zero. Indeed, otherwise, the restriction of the polynomial $f$ to this curve would be a polynomial of degree at most six with eight roots. Therefore, the sign of the coefficient $a$ does not depend on the choice of the four points inside the ovals and defines an invariant of the connected component of the discriminant complement. Any polynomial $f(x,y)$ in the considered component and its mirror image f(x,y)f(x,-y) define the opposite signs. The same is true for polynomials that realize the picture (e~)($\tilde${e}). $\Box$
@@ 3.3. Components for $J_{10}^{3}$ singularities
There are exactly 23 virtual components of the formal graph of type $J_{10}^{3}$. The shapes of the zero-level sets of some polynomials $f(x,y)$ representing thirteen of these components are shown in Fig. 5. In all these cases except those shown in pictures $(h)$, $(l)$, and $(m)$, the polynomials $f(x,y)$ and f(x,y)-f(-x,y), which are related by the involution ((34)), represent different virtual components. $\Box$
@@
whose zero-level set is shown in Fig. 6. Then, we add the term $\varepsilon y^{6}$ with sufficiently small $\varepsilon>0$ so that the principal weighted homogeneous part will be of class $J_{10}^{3}$. The obtained polynomial will then keep eight real critical points (three minima and five saddlepoints), which are visible in this figure, and do not significantly change their critical values. It will also have two critical points in the complex domain. Then we add a very small positive constant to the obtained polynomial in such a way that all saddlepoints will have positive critical values and all minima will stay in the negative domain. The zero-level set of the resulting polynomial will have the desired shape shown in picture (e) of Fig. 5. Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal weighted homogeneous part of the resulting polynomial will have the canonical form, see (32). Finally, we apply a parallel translation of the $x$ coordinate to kill the monomial $x^{2}$. The resulting polynomial realizes the picture $(e)$. It is invariant under the involution (33), in particular it has the form (32) with $\lambda_{2}=\lambda_{4}=\lambda_{7}=\lambda_{9}=0$. $\Box$
@@ 4. The case of $P_{8}$ singularities
4.2. The homological invariant of real and virtual components
@@
Consider three subsets in ${\mathbb{R}}^{3}$: $W_{0}=f^{-1}((-\infty,0])$, $W_{-}=f^{-1}((-\infty,T_{-}]),$ and $W_{+}=f^{-1}((-\infty,T_{+}])$. If $f\in\Xi_{1}$ (respectively, $f\in\Xi_{2}$) then the space $W_{-}$ is contractible (respectively, consists of two components, one homotopy equivalent to a point and the other to a circle). In particular, the reduced homology groups $\tilde{H}_{i}(W_{-})$ for $f\in\Xi_{2}$ are equal to $\mathbb{Z}$} for $i=0$ or 1 and are trivial for all other $i$.
@@
The homology index $\mbox{HI}(f)$ of a non-discriminant polynomial $f\in\Xi_{1}$ or $\Xi_{2}$ is the triple of Betti numbers $(b_{0},b_{1},b_{2})$ of the relative homology group $H_{*}(W_{0},W_{-})$.
@@ 4.3. The symmetry group and the relation between the real and virtual components
Proposition 33 (see [29], Corollary 8).
@@
If $f$ and $\tilde{f}$ are two functions related by the involution (45), then the corresponding exact sequences (43) are related by a kind of Poincaré duality: the group $H_{*}(W_{+},W_{0})$ for the polynomial $f$ is naturally dual to $H_{*}(W_{0},W_{-})$ for $\tilde{f}$, and vice versa.
@@
Proof. According to Lemma 35 of [29], there are only six polynomials of form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$. Three of these polynomials (mentioned in formula (26) of [29]) have signature (0,2) at the $A_{2}$ point. The remaining three polynomials have the form (46), in particular belong to the same orbit of the group $S(3)$. $\Box$
@@
The set of all polynomials of class $\Xi_{2}$ that have a critical point of class $E_{6}$ with zero critical value, and also two real critical points of class $A_{1}$ with the same critical value $\frac{64}{3}+\frac{128}{3\sqrt{3}}$, consists of three polynomials: the polynomial
@@
Proof of Proposition 34. The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 35 (respectively, 36) repeats the deduction of Proposition 9 from Lemma 10. Specifically, the neighborhood of the polynomial (46) or (47) in $\Xi_{2}$ has only one non-trivial automorphism that commutes with the Lyashko–Looijenga map: the involution $f(x,y,z)\leftrightarrow f(y,x,z)$. $\Box$
@@ 4.4. Components of discriminant complements of $P_{8}^{1}$ singularities
Each virtual component of type $P_{8}^{1}$ is associated with one and only one connected component of the space of non-discriminant polynomials of class $\Xi_{1}$. In particular, there are exactly seven such connected components.
@@
The two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2. For the first component, this invariant is equal to $(100)$, and for the second it is equal to $(001)$. According to the previous paragraph, the associated virtual components are also different.
@@
2. Let $\alpha$ be the coefficient at the monomial $q_{l}^{2}$ of the polynomial $\psi_{2}$ in (54). Then the lowest weighted homogeneous part of the polynomial $F(l)$ with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$. Substituting $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2}$ and $P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial into the polynomial with the lowest weighted homogeneous part
@@
The restriction of the polynomial $F(l)$ to the line connecting the origin and this Morse point is a degree three polynomial with vanishing derivative at these points; its critical point at the origin is a local maximum. Therefore, the critical value of $F(l)$ at the Morse critical point is negative. $\Box$
Denote by $C(l)$ the critical value of the polynomial $F(l)$ at its Morse critical point. Then the polynomial
$\tilde{F}(l)\equiv F(l)-\frac{C(l)}{2}$
is non-discriminant and has $\mbox{Ind}=-1$.
The polynomial $\tilde{F}(l)$ is canonically determined by the linear form $l$ and depends continuously on it. The union of all polynomials $\tilde{F}(l),$ $l\in L$, is connected and invariant under the action of the group $S(3)$. Therefore, the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement. $\Box$
@@ 4.5. One-dimensional cohomology of the discriminant complement of $P_{8}^{1}$ singularities
Proof. The zero-level sets of the polynomials in this component are homeomorphic to the plane with a handle. Indeed, the topology of these sets remains constant along the component. For any polynomial $\tilde{F}(l)$ from the previous proof this set is obtained from the very low level sets (homeomorphic to a plane) by passing through a Morse critical point of Morse index 1.
Thus, the one-dimensional homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$. These homology groups form a local system over the space of all such polynomials, i.e., over the connected component under consideration. In particular, the fundamental group of this component acts on these groups by monodromy. For each $l\in L$, one basic homology class of the zero-level set of the polynomial $\tilde{F}_{l}$ is defined by the intersection of this set and the plane $\ker l$. When $l\in L$ runs a half of the circle $L$, the plane $\ker l$ runs a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomial $\tilde{F}(l)$ runs a closed cycle in the discriminant complement. The monodromy over the latter cycle brings our basic 1-cycle to itself with the opposite orientation. Thus, the half-circle $L/{\mathbb{Z}}_{2}$ defines non-trivial elements of the fundamental group and of the mod 2 homology group of the component. $\Box$
@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
The invariant Card of a virtual component of the formal graph of an arbitrary singularity is the number of vertices in that virtual component.
@@
All of the polynomials given in the formulas (55), (56), (57), and (58) below are invariant under the $S(3)$-action. Thus, if $f$ is any of these polynomials and $C$ is a noncritical value of $f$, then Proposition 34 implies that the virtual component associated with the polynomial fCf-C is associated with only one real component of the discriminant complement. We will show that all virtual components with Card equal to 258, 336, 1648, 262, or 94 can be realized in this way. Hence, the assertion of Theorem 42 is true for these components.
Components with $\mbox{\rm Card}=1648$. For each polynomial $f$ of class $\Xi_{2}$, the polynomials fCf-C with sufficiently large $|C|$ and different signs of $C$ realize two different virtual components with $\mbox{Ind}=0$. The homology indices of these components are equal to $(000)$ and $(011)$. For example let $f$ be the $S(3)$-invariant polynomial
@@
has only two real Morse critical points: one with Morse index 2 and critical value 2-2, and the other with Morse index 1 and critical value 2. Thus, it has $\mbox{Ind}=1$ and homology index $(001)$. This polynomial is invariant under the involution (45). Therefore, according to the last statement of Proposition 41, the associated virtual component cannot have $\mbox{Card}=94$. Thus, this is the only remaining component with $\mbox{Ind}=1$, which has $\mbox{Card}=262$.
@@
has eight real Morse critical points: one local minimum with critical value 0, three critical points with Morse index $1$ and critical value $\frac{88}{49}$, three critical points with Morse index $2$ and critical value $4$, and one point with Morse index $1$ and critical value $12$. Therefore, the invariant $\mbox{Ind}(f-C)$ is equal to 0 for $C<0$, to $1$ for $C\in\left(0,\frac{88}{49}\right),$ to 2-2 for $C\in\left(\frac{88}{49},4\right),$ to 1 for $C\in\left(4,12\right),$ and to 0 for $C>12$.
The group $H_{*}(W_{0},W_{-})$ corresponding to any polynomial fCf-C with $C\in\left(0,\frac{88}{49}\right)$ is isomorphic to $\mathbb{Z}$} in dimension 0 and is trivial in all other dimensions. Therefore, the homology index $\mbox{HI}(f-1)$ is $(100)$. The involution (45) moves this polynomial to that with $\mbox{HI}=(012)$. In particular, these two polynomials represent two different components. They are $S(3)$-invariant, therefore by Proposition 34, the associated virtual components are also different. According to Proposition 41 and Table 3, these are two components with $\mbox{Card}=94$.
@@
Fig. 7. This graph does not contain edges of multiplicities other than 1 and 1-1. Upon a special request, our program verified that, for each virtual function of $P_{8}^{2}$ type from the virtual components with $\mbox{Card}=60$ or $156$, the intersection index of some two different vanishing cycles is equal to 2-2. Therefore, the virtual component associated with our polynomial is the only remaining virtual component with $\mbox{Ind}=-3$. $\Box$
@@
It is easy to check that the polynomials fCf-C with $\mbox{Ind}=-2$ and $\mbox{Ind}=1$ give new realizations of some components with $\mbox{Card}=336$ and $94$, respectively.
@@
Figure 8. Coxeter-Dynkin code for (59)
@@
According to Proposition 33, there exists a real polynomial $f_{0}\in\Xi_{2}$ with these data. According to Proposition 41 of [29], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), the odd Morse indices of their critical points are all equal to 1, and the even Morse indices are all equal to 2.
Let $i:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $i(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by increasing value, and $i(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$. Using the Lyashko–Looijenga covering (whose existence for the spaces $\Xi_{2}$ is shown in [12]) one can uniquely construct a path $I:[0,1)\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $i(t)$ is the sequence of critical values of the polynomial $I(t)$. As in the proof of Theorem 7 of [29], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined. Indeed, according to Proposition 2 of [12], this is guaranteed by the fact that all intersection forms corresponding to the limit positions $c_{-}$ and $c_{+}$ of the critical values of the polynomials $I(t)$ are elliptic. (These forms are given respectively by the upper left and the bottom right $3\times 3$ corners of the matrix (59).) The point $I(1)$ is a Morse (albeit non-generic) polynomial of class $\Xi_{2}$ having three critical points with critical value $c_{-}$ and three critical points with critical value $c_{+}$.
@@
The involution (45) moves the polynomial $I(1)$ to a polynomial in the other component of the discriminant complement with $\mbox{Card}=60$. Thus, this component is also represented by an $S(3)$-invariant polynomial. This proves the assertion of Theorem 42 for both virtual components with $\mbox{Card}=60$.
@@
  • its critical point of class $D_{4}^{-}$ splits into one minimum point with critical value $\varepsilon$, and three critical points with Morse index 1 and critical values 0, $\varepsilon^{2},$ and $2\varepsilon^{2}$,

### v5

@@ 2.5. The components for the $X_{9}^{1}$ singularities
Each polynomial (15) is invariant under the transformation $(x,y)\mapsto(y,x)$. The transformation $(x,y)\mapsto(-x,-y)$ maps each such polynomial to the analogous polynomial with the opposite value of parameter $t$. In particular, it permutes the connected components of the discriminant complement that are described by pictures $(d)$ and $(e)$.
@@ 2.6. Components for the $X_{9}^{2}$ singularities
The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the discriminant complement in some way. All these components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
@@
Each polynomial of degree four on ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by an affine transformation of ${\mathbb{R}}^{2}$ to the form
@@
Proof of Lemma 23. This lemma follows from Proposition 22 of [28] by the substitution $\tilde{x}=|a|^{\frac{3}{4}}x$, $\tilde{y}=\mbox{sign}(a)|a|^{\frac{-1}{4}}y$, $t=|a|^{\frac{-5}{4}}$, and the sign $\pm$ in (18) equal to $\mbox{sign}(a)$. $\Box$
@@ 3. Components of discriminant complements for $J_{10}$ singularities
Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (31) and (32), respectively. The virtual functions associated with generic polynomials of these types, the formal graphs of types $J_{10}^{1}$ and $J_{10}^{3}$, and their virtual components are defined in the same way as for the case $X_{9}$, except that the sets of vanishing cycles consist of ten elements.
@@
In the following pictures (4)–(6), it is assumed that the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
@@ 3.2. Components for $J_{10}^{1}$ singularities
The following is an invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(e)$ from its mirror image f(x,y)f(x,-y). The four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement. Let us take an arbitrary point inside each of these four ovals. There is a single curve with the equation $x=ay^{3}+by^{2}+cy+d$ that passes through these four points. The coefficient $a$ of this equation cannot be zero. Otherwise, the restriction of the polynomial $f$ to this curve would be a polynomial in the coordinate $y$ of degree at most six with eight roots. Therefore, the sign of the coefficient $a$ does not depend on the choice of the four points inside the ovals and defines an invariant of the connected component of the discriminant complement. Any polynomial $f(x,y)$ in the considered component and its mirror image f(x,y)f(x,-y) define the opposite signs. The same is true for polynomials that realize the picture (e~)($\tilde${e}). $\Box$
@@ 3.3. Components for $J_{10}^{3}$ singularities
There are exactly 23 virtual components of the formal graph of type $J_{10}^{3}$. The shapes of the zero-level sets of some polynomials $f(x,y)$ representing thirteen of these components are shown in Fig. 5. In all these cases except those shown in pictures $(g)$, $(l)$, and $(m)$, the polynomials $f(x,y)$ and f(x,y)-f(-x,y), which are related by the involution ((34)), represent different virtual components. $\Box$
@@
whose zero-level set is shown in Fig. 6. Then, we add the term $\varepsilon y^{6}$ with sufficiently small $\varepsilon>0$ so that the principal weighted homogeneous part will be of class $J_{10}^{3}$. The obtained polynomial will then keep eight real critical points (three minima and five saddlepoints), which are visible in this figure, and do not significantly change their critical values. It will also have two critical points in the complex domain. Then we add a very small positive constant to the obtained polynomial in such a way that all saddlepoints will have positive critical values and all minima will stay in the negative domain. The zero-level set of the resulting polynomial will have the desired shape shown in picture $(e)$ of Fig. 5, see Fig. 6 right. Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal weighted homogeneous part of the resulting polynomial will have the canonical form, see (32). Finally, we apply a parallel translation of the $x$ coordinate to kill the monomial $x^{2}$. The resulting polynomial realizes the picture $(e)$. It is invariant under the involution (33), in particular it has the form (32) with $\lambda_{2}=\lambda_{4}=\lambda_{7}=\lambda_{9}=0$. $\Box$
@@ 4. The case of $P_{8}$ singularities
4.2. The homological invariant of real components
@@
Consider three subsets in ${\mathbb{R}}^{3}$: $W_{0}=f^{-1}((-\infty,0])$, $W_{-}=f^{-1}((-\infty,T_{-}]),$ and $W_{+}=f^{-1}((-\infty,T_{+}])$. If $f\in\Xi_{1}$ (respectively, $f\in\Xi_{2}$), then the space $W_{-}$ is contractible (respectively, consists of two components, one homotopy equivalent to a point and the other to a circle). In particular, the reduced homology groups $\tilde{H}_{i}(W_{-})$ for $f\in\Xi_{2}$ are equal to $\mathbb{Z}$} for $i=0$ or 1 and are trivial for all other $i$.
@@
The homology index $\mbox{HI}(f)$ of a non-discriminant polynomial $f$ of class $\Xi_{1}$ or $\Xi_{2}$ is the triple of Betti numbers $(b_{0},b_{1},b_{2})$ of the relative homology group $H_{*}(W_{0},W_{-})$.
@@ 4.3. The symmetry group and the relation between the real and virtual components
Proposition 33 (see [29], Theorem 10).
For any generic Morse polynomial $f$ of type $\Xi_{1}$ or $\Xi_{2}$, a virtual Morse function $\varphi$ associated with it, and a virtual Morse function $\tilde{\varphi}$ connected with $\varphi$ by an edge of the formal graph, there exists a generic Morse polynomial $\tilde{f}$ and a path in the space $\Xi_{1}$ or $\Xi_{2}$ connecting $f$ and $\tilde{f}$ and containing only one non-generic point at which it experiences a standard surgery of the same type as the edge $\varphi,\tilde{\varphi}$].
Corollary 34.
@@
If $f$ and $\tilde{f}$ are two polynomials related by the involution (45), then the corresponding exact sequences (43) are related by a kind of Poincaré duality: the group $H_{*}(W_{+},W_{0})$ for the polynomial $f$ is naturally dual to $H_{*}(W_{0},W_{-})$ for $\tilde{f}$, and vice versa.
@@
Proof. According to Lemma 38 of [29], there are only six polynomials of the form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$. Three of these polynomials (mentioned in formula (27) of [29] with a certain value of the parameter $\beta$) have signature (0,2) at the $A_{2}$ point. The remaining three polynomials have the form (46) up to the permutations of coordinates, see formula (28) in [29]. In particular, they belong to one orbit of the group $S(3)$. $\Box$
@@
The set of all polynomials of of the form (42) and class $\Xi_{2}$ that have a critical point of class $E_{6}$ with zero critical value, and also two real critical points of class $A_{1}$ with the same critical value $\frac{64}{3}+\frac{128}{3\sqrt{3}}$, consists of three polynomials: the polynomial
@@
Proof of Proposition 35. The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 36 (respectively, 37) repeats the deduction of Proposition 9 from Lemma 10. Specifically, the neighborhood of the polynomial (46) or (47) in $\Xi_{2}$ has only one non-trivial automorphism that commutes with the Lyashko–Looijenga map: the involution $f(x,y,z)\leftrightarrow f(y,x,z)$. $\Box$
@@ 4.4. Components of discriminant complements of $P_{8}^{1}$ singularities
Each virtual component of type $P_{8}^{1}$ is associated with one and only one connected component of the space of non-discriminant polynomials in class $\Xi_{1}$. In particular, there are exactly seven such connected components.
@@
The two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2. For the first component, this invariant is equal to $(100)$, and for the second it is equal to $(001)$. According to the last sentence of the previous paragraph, the associated virtual components are also different.
@@
2. Let $\alpha$ be the coefficient at the monomial $q_{l}^{2}$ of the polynomial $\psi_{2}$ in (54). Then the lowest weighted homogeneous part of the polynomial (52) with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$. Substituting $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2}$ and $P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial into the polynomial with the lowest weighted homogeneous part
@@
The restriction of the polynomial (52) to the line connecting the origin and this Morse point is a degree three polynomial with vanishing derivative at these points; its critical point at the origin is a local maximum. Therefore, its critical value at the Morse critical point is negative. $\Box$
Denote by $C(l)$ the critical value of the polynomial (52) at its Morse critical point. Let $F(l)$ be the polynomial obtained from (52) by adding the constant $C(l)/2$. Then the polynomial $F(l)$ is non-discriminant and has $\mbox{Ind}=-1$.
The latter polynomial is canonically determined by the linear form $l$ and depends continuously on it. The union of polynomials $F(l)$ over all $l\in L$ is connected and invariant under the action of the group $S(3)$. Therefore, the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement. $\Box$
@@ 4.5. One-dimensional cohomology of the discriminant complement of $P_{8}^{1}$ singularities
Proof. The zero-level sets of the polynomials in this component are homeomorphic to the plane with a handle. Indeed, the topology of these sets remains constant along the component. For any polynomial $F(l)$ from the previous proof this set is obtained from the very low level sets (homeomorphic to a plane) by passing through a Morse critical point of Morse index 1.
Thus, the one-dimensional homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$. These homology groups form a local system over the space of all such polynomials, i.e., over the connected component of the discriminant complement under consideration. In particular, the fundamental group of this component acts on these groups by monodromy. For each $l\in L$, one basic homology class of the zero-level set of the polynomial $F_{l}$ is defined by the intersection of this set and the plane $\ker l$. When $l\in L$ runs a half of the circle $L$, the planes $\ker l$ run a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomials $F(l)$ run a closed cycle in the discriminant complement. The monodromy over the latter cycle brings our basic 1-cycle to itself with the opposite orientation. Thus, the half-circle $L/{\mathbb{Z}}_{2}$ defines non-trivial elements of the fundamental group and of the mod 2 homology group of the component. $\Box$
@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
The invariant Card of a virtual component of the formal graph is the number of vertices in that virtual component.
@@
All of the polynomials given in the formulas (55), (56), (57), and (58) below are invariant under the $S(3)$-action. Thus, if $f$ is any of these polynomials and $C$ is a noncritical value of $f$, then Proposition 35 implies that the virtual component associated with the polynomial fCf-C is associated with only one real component of the discriminant complement. We will show that all virtual components with Card equal to 258, 336, 1648, 262, or 94 can be realized in this way. Hence, the assertion of Theorem 43 is true for these seven components.
Components with $\mbox{\rm Card}=1648$. For each polynomial $f$ of class $\Xi_{2}$, the polynomials fCf-C with sufficiently large $|C|$ and different signs of $C$ realize two different virtual components with $\mbox{Ind}=0$. The homology indices of these components are equal to $(000)$ and $(011)$. For example, let $f$ be the $S(3)$-invariant polynomial
@@
has only two real Morse critical points: one with Morse index 2 and critical value 2-2, and the other with Morse index 1 and critical value 2. Thus, it has $\mbox{Ind}=1$ and homology index $(001)$. This polynomial is invariant under the involution (45). According to the last statement of Proposition 42, the associated virtual component cannot have $\mbox{Card}=94$. Therefore, this is the only remaining component with $\mbox{Ind}=1$, which has $\mbox{Card}=262$.
@@
has eight real Morse critical points: one local minimum with critical value 0, three critical points with Morse index $1$ and critical value $\frac{88}{49}$, three critical points with Morse index $2$ and critical value $4$, and one critical point with Morse index $1$ and critical value $12$. Therefore, the invariant $\mbox{Ind}(f-C)$ is equal to 0 for $C<0$, to $1$ for $C\in\left(0,\frac{88}{49}\right),$ to 2-2 for $C\in\left(\frac{88}{49},4\right),$ to 1 for $C\in\left(4,12\right),$ and to 0 for $C>12$.
The group $H_{*}(W_{0},W_{-})$ corresponding to any polynomial fCf-C with $C\in\left(0,\frac{88}{49}\right)$ is isomorphic to $\mathbb{Z}$} in dimension 0 and is trivial in all other dimensions. Therefore, the homology index $\mbox{HI}(f-1)$ is $(100)$. The involution (45) moves this polynomial to one with $\mbox{HI}=(012)$. In particular, these two polynomials represent two different components. They are $S(3)$-invariant. Therefore, by Proposition 35, the associated virtual components are also different. According to Proposition 42 and Table 3, these are two components with $\mbox{Card}=94$.
@@
Fig. 7. This graph does not contain edges of multiplicities other than 1 and 1-1. Upon a special request, our program verified that, for each virtual function of $P_{8}^{2}$ type from the virtual components with $\mbox{Card}=60$ or $156$, the intersection index of some two different vanishing cycles of the canonical basis is equal to 2-2. Therefore, the virtual component associated with our polynomial is the only remaining virtual component with $\mbox{Ind}=-3$. $\Box$
@@
It is easy to check that the polynomials fCf-C with $f$ given by (58) and $C\in\left(\frac{88}{147},\frac{4}{3}\right)$ or $C\in\left(0,\frac{88}{147}\right)$ give new realizations of components with $\mbox{Card}=336$ and $94$, respectively.
@@
Figure 8. Coxeter-Dynkin graph for (59)
@@
According to Corollary 34, there exists a real polynomial $f_{0}\in\Xi_{2}$ with these data. According to Proposition 41 of [29], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), the odd Morse indices of their critical points are all equal to 1, and the even Morse indices are all equal to 2. Therefore the polynomial $f_{0}$ has no local minima or maxima.
Let $s:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $s(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by increasing value, and $s(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$. Using the Lyashko–Looijenga covering (whose existence for the spaces $\Xi_{2}$ is shown in [12]) one can uniquely construct a path $I:[0,1)\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $s(t)$ is the sequence of critical values of the polynomial $I(t)$. As in the proof of Theorem 10 of [29], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined. Indeed, according to Proposition 2 of [12], this is guaranteed by the fact that all intersection forms corresponding to the limit positions $c_{-}$ and $c_{+}$ of the critical values of the polynomials $I(t)$ are elliptic. (These forms are given respectively by the upper left and the bottom right $3\times 3$ corners of the matrix (59).) The point $I(1)$ is a Morse (albeit non-generic) polynomial of class $\Xi_{2}$ having three critical points with critical value $c_{-}$ and three critical points with critical value $c_{+}$.
@@
By statement 3 of Proposition 42, the involution (45) moves the polynomial $I(1)$ to a polynomial in the different component of the discriminant complement with $\mbox{Card}=60$. Thus, this component is also represented by an $S(3)$-invariant polynomial. This proves the assertion of Theorem 43 for both virtual components with $\mbox{Card}=60$.
@@
  • its critical point of class $D_{4}^{-}$ splits into one minimum point with critical value $\varepsilon$, and three critical points with Morse index 1 and critical values 0, $\varepsilon^{2},$ and $\varepsilon^{2}$,